# Geometric Bounds for Persistence

### v1

@@ 5 Bounding Extinction Times
In Section 3, we discussed bounds for the lifespans of individual homology classes in degree $\geq k$ via different notions of $k$-width from metric geometry. However, as Figure 2 suggests, these $k$-widths do not give any bound for the lower homology classes in lower degrees. In other words, if $\operatorname{AW}_{k}(\mathcal{M}\subset\mathbb{R}^{N})$ (or another $k$-width) is small, the filling radius (or lifespan) of a homology class $\alpha\in\operatorname{\tilde{H}}_{j}(\mathcal{M})$ for $j<k$ can still be very large. It is easy to see that the example in Figure 2 can be generalized to any dimension and codimension. In particular, $k$-widths do not say much about the size of topological features in dimensions lower than $k$.
In this section, we use a different approach and propose a global bound to the lifespans of homology classes in all degrees at once. Our goal is to bound the following quantity:
@@
$\xi(\mathcal{X})=\sup_{k\geq 0}\sup\{d_{\omega}\mid\omega\in\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))\}$
is called the VR-extinction time of $\mathcal{X}$. Similarly, define $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ as the Čech-extinction time for Čech filtration in a Banach space $\mathbf{Y}$. We will use the terms VR-extinction (respectively Čech-extinction) for short.
Notice that with the VR-extinction $\xi(\mathcal{X})$ time, we are not only bounding the lifespans of homological features with birth time $=0$, but we are bounding the death time of all homological features (in degree $1$ and higher) appearing throughout the filtration. Also, note that the radius of $\mathcal{X}$ (2.35) automatically gives an upper bound for VR-extinction time, i.e. $\xi(\mathcal{X})\leq\mathrm{rad}(\mathcal{X})$ by [LMO23, Remark 9.4]). A similar statement is clearly true for Čech-extinction time. i.e., $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})\leq\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$.
Remark 5.2.
Due to 2.21, any bound on extinction times of a filtration will result in an upper bound on the right endpoints of any interval in the barcode of said filtration; cf. 2.22.
@@
$\mathrm{E}=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{N}\dfrac{x_{i}^{2}}{a_{i}^{2}}=1\right\},$
@@
$\overline{\mathrm{E}}=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{N}\dfrac{x_{i}^{2}}{a_{i}^{2}}\leq 1\right\},$
be the solid ellipsoid. Given an integer $m\geq 1$, we will use to notation $\overline{\mathrm{E}}^{m}=\{\mathbf{x}\in\overline{\mathrm{E}}\mid x_{i}=0\mbox{ for }i>m\}$. Let $\{D_{1},D_{2},\dots,D_{\ell_{k}}\}$ be $\ell_{k}$ disjoint small disks in $\overline{\mathrm{E}}^{N-k}$. Then, consider the following surgery operation. For $1\leq j\leq\ell_{k}$, let $\Omega_{j}=\overline{\mathrm{E}}\cap(D_{j}\times\mathbb{R}^{k})$. Let $S_{j}=\mathrm{E}\cap\Omega_{j}$, and $T_{j}=\operatorname{cl}(\partial\Omega_{j}\setminus S_{j})$. Then, swapping the $S_{j}$ and the $T_{j}$ will give a new closed manifold
$\widehat{\mathrm{E}}=\left(\mathrm{E}\setminus\bigcup_{j=1}^{\ell_{k}}S_{j}\right)\cup\bigcup_{j=1}^{\ell_{k}}T_{j}\subset\mathbb{R}^{N}.$
See Figure 4 for an illustration. While $\operatorname{\tilde{H}}_{k}(\mathrm{E})$ is trivial for $10<k<N-$ 1, $\mbox{rank}(\operatorname{\tilde{H}}_{k}(\widehat{\mathrm{E}}))\geq{\ell_{k}}$ because of the homology classes generated by the $k$-handles $\{T_{j}\}$. By choosing $\{D_{j}\}$ such that $\{\Omega_{j}\}$ are all pairwise disjoint, one can obtain $\widehat{\mathrm{E}}$ with nontrivial homology in all desired degrees $k$.
@@
Notice that, through Kolmogorov widths, Section 1 yields that
@@
for any class $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathrm{R}^{N})).$ By suitably choosing the numbers $\{a_{i}\}$, all $k$-widths for $k\leq\dim\widehat{\mathrm{E}}-1$ can be made arbitrarily large. However, it is not hard to see that, in this particular example, the death times of all these homology classes do not exceed the smallest axis length $a_{N}$. This example therefore shows that $k$-widths can highly overestimate lifespans. In other words, we have that the Čech extinction satisfies $\widecheck{\xi}(\widehat{\mathrm{E}}\subset\mathbb{R}^{N})\leq a_{N}$. In the following sections, we estimate extinction radii by comparing a given space to a nearby topologically trivial space.
5.1 Bounding Čech-extinction with Convex Hulls
@@
$\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y})=\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathrm{conv}(\mathcal{X}))=\sup_{y\in\mathrm{conv}(\mathcal{X})}\inf_{x\in\mathcal{X}}\|y-x\|.$
@@
The stability theorems (Section 2.1.3) do not imply the previous result as they would only yield bounds on lifespans but not on death times. Indeed, 2.16 implies that
@@
Since $\widecheck{\mathrm{PD}}_{k}(\mathrm{conv}(\mathcal{X})\subset\mathbf{Y})=\emptyset$, the stability result and the definition of the bottleneck distance [EH10] imply that for any $(b,d)\in\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$,
@@
However, both $b$ and $d$ can be arbitrarily large. Meanwhile, Section 1 bounds elements of $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ independently of $k$: for any $(b,d)\in\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$,
@@
Furthermore:
  • Section 1 applies to all elements of $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ – not just those that can be associated to an interval in $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$; see 5.2.
  • The stability theorems do not in general apply in order to obtain bounds for either the lifetime or extinction of arbitrary classes in $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ or $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. For example, the Filling Radius of a manifold is not stable under the Gromov-Hausdorff distance; see [LMO23, Sections 9.4 and 9.5].
@@
Example in Figure 3 shows that it is not true that the extinction $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ is upper-bounded by the überspread $\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y})$.
While the above remark shows that the distance to the nearest übercontractible space (überspread) fails to bound extinction, it might be still true that distance to a specific cleverly chosen übercontractible space might give an estimate on extinction. One natural choice for the role of an übercontractible space $\mathcal{T}$ approximating $\mathcal{X}\subset\mathbf{Y}$ is the convex hull of $\mathcal{X}$; this is exactly what has just been discussed; with this choice, there is an extinction bound (Theorem 1), but it is tempting to improve it by choosing a finer $\mathcal{T}$. Another natural choice of $\mathcal{T}$ is given by the cut-locus of the boundary of the convex hull of $\mathcal{X}$ (see the discussion in Appendix B). Unfortunately, the extinction time is not bounded from above by the distance from $\mathcal{X}$ to this $\mathcal{T}$ (chosen as the cut-locus of $\partial\mathrm{conv}(\mathcal{X})$, assuming $\mathrm{conv}(\mathcal{X})$ full-dimensional). We omit the discussion of examples in view of the negative nature of the result, but they can be obtained as subsets of a square (one can take a square and cut out a large off-centered disk).
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Let $\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})}$ be the cut-locus of the convex hull of $\mathcal{X}$ (see Appendix B). Is it possible to upper-bound the extinction $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ in terms of the Hausdorff distance to $\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})}$? i.e. for some $C_{\mathbf{Y}}>0$, is it true that
@@ 5.2 Bounding VR-extinction via the Tight Span
To get a bound on VR-extinctions, Theorem 1 can be applied for the Kuratowski distance-preserving embedding of $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$, but the result will be sub-optimal. We refine it now.
The following definition is analogous to Definition 5.4 in that the convex hull of $\mathcal{X}\subset\mathbf{Y}$ is supplanted by $\operatorname{E}(\mathcal{X})$ (see 2.27). Recall that $\mathcal{X}$ naturally embeds in its tight span, so we can assume that $\mathcal{X}\subset\operatorname{E}(\mathcal{X})$.
@@
$\operatorname{hcdef}(\mathcal{X}):=\mathbf{d}_{\mathrm{H}}^{\operatorname{E}(\mathcal{X})}(\mathcal{X},\operatorname{E}(\mathcal{X})).$
The following corollary to 2.7 is analogous to Theorem 1.
@@
Let $\mathcal{X}$ be the unit $\ell^{\infty}$ sphere in $\mathbb{R}^{2}$. Then, in that case, by results of [KK16], $\operatorname{E}(\mathcal{X})\cong\Big{(}[-1,1]\times[-1,1],\ell^{\infty}\Big{)}$ and we compute that $\operatorname{hcdef}(\mathcal{X})=1$. Since $\mathcal{N}_{t}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\simeq\mathcal{X}\simeq S^{1}$ for every $10<t<$ 1, by Theorem 2.7 we see that, for $\omega=[S^{1}]$, $d_{\omega}=2$ (see also [LMO23, Corollary 7.4].
Recall from [LMO23, Proposition 9.1] that $\mathcal{V}_{r}(\mathcal{X})$ becomes contractible as soon as $r\geq\mathrm{rad}(\mathcal{X})$. The next proposition proves that the bound in Corollary 5.9 is never worse than this bound.
@@
$\operatorname{hcdef}(\mathcal{X})\leq\mathrm{rad}(\mathcal{X})$
hold for every compact metric space $\mathcal{X}$.
@@
Assume that $\delta>\mathrm{rad}(\mathcal{X})$ and let $x_{0}\in\mathcal{X}$ be a point such that $\mathbf{d}_{\mathcal{X}}(x_{0},x)<\delta$ for all $x\in\mathcal{X}$. Recall that, according to equation (4), we have that
@@
Let $x^{\prime}_{0}\in\mathcal{X}$ be such that $f(x_{0})=\mathbf{d}_{\mathcal{X}}(x_{0},x^{\prime}_{0})-f(x_{0}^{\prime})$. Notice that then, for a given for $f\in\operatorname{E}(\mathcal{X})$, by equation (4) we have both
@@
Adding these two expressions we obtain that
@@
$\min\big{(}\|f-\mathbf{d}_{\mathcal{X}}(x_{0},\cdot)\|_{\infty},\|f-\mathbf{d}_{\mathcal{X}}(x_{0}^{\prime},\cdot)\|_{\infty}\big{)}\leq\frac{\delta}{2}.$
Since $f\in\operatorname{E}(\mathcal{X})$ was arbitrary, this establishes that $\operatorname{hcdef}(\mathcal{X})\leq\frac{\delta}{2}$ from which the claim follows. ∎

### v2

@@ 5 Bounding Extinction Times
In Section 3, we discussed bounds for the lifespans of individual homology classes in degree $\geq k$ via different notions of $k$-width from metric geometry. However, as Figure 2 suggests, these $k$-widths do not give any bound for the lower homology classes in lower degrees. In other words, if $\operatorname{AW}_{k}(\mathcal{M}\subset\mathbb{R}^{N})$ (or another $k$-width) is small, the filling radius (or lifespan) of a homology class $\alpha\in\widetilde{\mathrm{H}}_{j}(\mathcal{M})$ for $j<k$ can still be very large. It is easy to see that the example in Figure 2 can be generalized to any dimension and codimension. In particular, $k$-widths do not say much about the size of topological features in dimensions lower than $k$.
In this section, we use a different approach and propose a global bound to the lifespans of homology classes in all degrees at once. Our goal is to bound the following quantity.
@@
$\xi(\mathcal{X}):=\sup_{k\geq 0}\sup\big{\{}d_{\omega}\mid\omega\in\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))\big{\}}$
is called the VR-extinction time of $\mathcal{X}$. We similarly, define $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$, the Čech-extinction time for the Čech filtration of a subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$. We will use the terms VR-extinction (respectively Čech-extinction) for short.
Notice that via the VR-extinction $\xi(\mathcal{X})$, we are not only bounding the lifespans of homological features with birth time $=0$, but we are bounding the death time of all homological features (in degree $1$ and higher) appearing throughout the filtration. Also, note that, as pointed out in [LMO24, Remark 9.8], the radius of $\mathcal{X}$ (2.39) automatically gives an upper bound for the VR-extinction time, that is, $\xi(\mathcal{X})\leq\mathrm{rad}(\mathcal{X})$. This is easy to see since the VR complex $\mathcal{V}_{r}(\mathcal{X})$ is a simplicial cone for $r\geq\mathrm{rad}(\mathcal{X})$. A similar statement is clearly true for Čech-extinction, i.e. $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})\leq\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$. See 5.2 for comments on the usefulness of this type of bounds.
Remark 5.2 (Utility of Extinction Time Bounds).
Bounds on extinction times can be useful in practice for reducing the computational effort incurred by algorithms designed for calculating PH, as we now explain. By 2.22, any bound on the extinction times of a compact metric space provides an upper bound on the right endpoints of every interval in the barcode of the VR-filtration for that space; see 2.23. This is particularly useful due to the inherent structure of algorithmic procedures for computing persistence diagrams of simplicial filtrations whose complexity increases with the total number of simplices; see [EH10, Chapter VII]. For instance, both the software packages Ripser [Bau21] and Eirene [HG16] use $\mathrm{rad}(X)$ as a cut-off value for the filtration parameter. Any computationally feasible approximation to the extinction time bounds below could similarly boost efficiency in practical applications.
@@
$\mathrm{E}:=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{N}\dfrac{x_{i}^{2}}{a_{i}^{2}}=1\right\},$
@@
$\overline{\mathrm{E}}:=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{N}\dfrac{x_{i}^{2}}{a_{i}^{2}}\leq 1\right\},$
be the solid ellipsoid. Given an integer $m\geq 1$, we will use to notation $\overline{\mathrm{E}}^{m}:=\{\mathbf{x}\in\overline{\mathrm{E}}\mid x_{i}=0\mbox{ for }i>m\}$. Let $\{D_{1},D_{2},\dots,D_{\ell_{k}}\}$ be $\ell_{k}$ disjoint small disks in $\overline{\mathrm{E}}^{N-k}$. Then, consider the following surgery operation. For $1\leq j\leq\ell_{k}$, let $\Omega_{j}:=\overline{\mathrm{E}}\cap(D_{j}\times\mathbb{R}^{k})$. Let $S_{j}:=\mathrm{E}\cap\Omega_{j}$, and $T_{j}:=\operatorname{cl}(\partial\Omega_{j}\setminus S_{j})$. Then, swapping the $S_{j}$ and the $T_{j}$ will give a new closed manifold
$\widehat{\mathrm{E}}:=\left(\mathrm{E}\setminus\bigcup_{j=1}^{\ell_{k}}S_{j}\right)\cup\bigcup_{j=1}^{\ell_{k}}T_{j}\subset\mathbb{R}^{N}.$
See Figure 4 for an illustration. While $\widetilde{\mathrm{H}}_{k}(\mathrm{E})$ is trivial for $10<k<N-$ 1, $\mbox{rank}(\widetilde{\mathrm{H}}_{k}(\widehat{\mathrm{E}}))\geq{\ell_{k}}$ because of the homology classes generated by the $k$-handles $\{T_{j}\}$. By choosing $\{D_{j}\}$ such that $\{\Omega_{j}\}$ are all pairwise disjoint, one can obtain $\widehat{\mathrm{E}}$ with nontrivial homology in all desired degrees $k$.
@@
Notice that, through Kolmogorov widths, Section 1.2 yields that
@@
for any class $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathrm{R}^{N})).$ By suitably choosing the numbers $\{a_{i}\}$, all $k$-widths for $k\leq\dim(\widehat{\mathrm{E}})-1$ can be made arbitrarily large. However, it is not hard to see that, in this particular example, the death times of all these homology classes do not exceed the smallest axis length $a_{N}$. This example therefore shows that $k$-widths can highly overestimate lifespans. In other words, we have that the Čech extinction satisfies $\widecheck{\xi}(\widehat{\mathrm{E}}\subset\mathbb{R}^{N})\leq a_{N}$. In the following sections, we estimate extinction radii by comparing a given space to a nearby topologically trivial space.
5.1 Bounding Čech-extinction via Convex Hulls
@@
$\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y}):=\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathrm{conv}(\mathcal{X}))=\sup_{y\in\mathrm{conv}(\mathcal{X})}\inf_{x\in\mathcal{X}}\|y-x\|.$
@@
The stability theorems (Section 2.1.3) do not imply the previous result for the following reasons:
  1. 1.
    In contrast with the stability theorems 2.15 and 2.16, which apply only to homology classes associated to points in the persistence diagram $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$, Section 1.2 applies to all elements of $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. Recall, from 2.21 and 2.22, that in general $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ contains strictly more homology classes than those which can be associated with points in the persistence diagram $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$.
  2. 2.
    Strictly speaking, one should not expect that, in general, stability holds for the lifetime or extinction of arbitrary classes in $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ or $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. For example, the filling radius of an $m$-dimensional manifold $\mathcal{M}$ coincides with both the lifetime and extinction of its fundamental class $[\mathcal{M}]\in\mathrm{Spec}_{m}(\mathcal{V}_{\bullet}(\mathcal{M}))$. However, the filling radius of $\mathcal{M}$ is not stable under the Gromov–Hausdorff distance as explained in [LMO24, Sections 9.4 and 9.5].
  3. 3.
    Even if we restricted ourselves to those homology classes associated with points in the persistence diagram, the stability theorems would only yield bounds on lifespans but not on death times (i.e. on extinction). Indeed, 2.16 implies that
@@
    Since $\widecheck{\mathrm{PD}}_{k}(\mathrm{conv}(\mathcal{X})\subset\mathbf{Y})=\emptyset$, the stability result and the definition of the bottleneck distance [EH10] imply that for any point $(b,d)\in\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$,
@@
    However, both $b$ and $d$ can be arbitrarily large without violating this inequality. Meanwhile, Section 1.2 bounds from above the second coordinate of every element of $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$: for any $(b,d)\in\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$,
@@
Example in Figure 3 shows that it is not true that the extinction $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ is bounded from above by the überspread $\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y})$.
While the above remark shows that the distance to the nearest übercontractible space (überspread) fails to bound extinction, it might be still true that distance to a specific cleverly chosen übercontractible space might give an estimate on extinction. One natural choice for the role of an übercontractible space $\mathcal{T}$ approximating $\mathcal{X}\subset\mathbf{Y}$ is the convex hull of $\mathcal{X}$; this is exactly what has just been discussed; with this choice, there is an extinction bound (Theorem 1.2), but it is tempting to improve it by choosing a finer $\mathcal{T}$. Another natural choice of $\mathcal{T}$ is given by the cut-locus of the boundary of the convex hull of $\mathcal{X}$ (see the discussion in Appendix B). Unfortunately, the extinction time is not bounded from above by the distance from $\mathcal{X}$ to this $\mathcal{T}$ (chosen as the cut-locus of $\partial\mathrm{conv}(\mathcal{X})$, assuming $\mathrm{conv}(\mathcal{X})$ full-dimensional). We omit the discussion of examples in view of the negative nature of the result, but they can be obtained as subsets of a square (one can take a square and cut out a large off-centered disk).
@@
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Let $\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})}$ be the cut-locus of the convex hull of $\mathcal{X}$ (see Appendix B). Is it possible to upper-bound the extinction $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ in terms of the Hausdorff distance to $\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})}$? That is, is it true that
@@ 5.2 Bounding VR-extinction via the Tight Span
To derive a bound on VR-extinction times, Theorem 1.2 can be applied to the Kuratowski distance-preserving embedding of $\mathcal{X}$ into $L^{\infty}(\mathcal{X})$. However, this approach yields a sub-optimal result. Here, we present a refinement.
The following definition is analogous to Definition 5.4 in that the convex hull of $\mathcal{X}\subset\mathbf{Y}$ is supplanted by $\operatorname{E}(\mathcal{X})$ (see 2.28). Recall that $\mathcal{X}$ naturally embeds in its tight span, so we can assume that $\mathcal{X}\subset\operatorname{E}(\mathcal{X})$.
@@
$\operatorname{hcdef}(\mathcal{X}):=\mathbf{d}_{\mathrm{H}}^{\operatorname{E}(\mathcal{X})}(\mathcal{X},\operatorname{E}(\mathcal{X}))=\sup_{f\in\operatorname{E}(\mathcal{X})}\inf_{x\in\mathcal{X}}\|f-\mathbf{d}_{\mathcal{X}}(x,\cdot)\|_{\infty}.$
The following corollary to 2.7 is analogous to Theorem 1.2.
@@
Let $\mathcal{X}$ be the unit $\ell^{\infty}$ sphere in $\mathbb{R}^{2}$. Then, in that case, by results of Kılıç and Koçak [KK16], $\operatorname{E}(\mathcal{X})$ is isometric to $\Big{(}[-1,1]\times[-1,1],\ell^{\infty}\Big{)}$ and we compute that $\operatorname{hcdef}(\mathcal{X})=1$. Since $\mathcal{N}_{t}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\simeq\mathcal{X}\simeq S^{1}$ for every $10<t<$ 1, and for $t\geq 1$ we have $\mathcal{N}_{t}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))=\operatorname{E}(\mathcal{X})$, which is contractible, by 2.7 we have
  • $\mathcal{V}_{r}(\mathcal{X})\simeq S^{1}$ for every $20<r<$ 2, and
  • $\mathcal{V}_{r}(\mathcal{X})$ is contractible for $r\geq 2$.
Hence, for $\omega=[S^{1}]$, $d_{\omega}=2$; see also [LMO24, Corollary 7.13].
Recall from [LMO24, Remark 9.8] that $\mathcal{V}_{r}(\mathcal{X})$ becomes contractible as soon as $r\geq\mathrm{rad}(\mathcal{X})$. The next proposition proves that the bound in Corollary 5.9 is never worse than this bound.
@@
$2\operatorname{hcdef}(\mathcal{X})\leq\mathrm{rad}(\mathcal{X})$
holds for every compact metric space $\mathcal{X}$.
@@
Assume that $\delta>\mathrm{rad}(\mathcal{X})$ and let $x_{0}\in\mathcal{X}$ be a point such that $\mathbf{d}_{\mathcal{X}}(x_{0},x)<\delta$ for all $x\in\mathcal{X}$. Pick any $f\in\operatorname{E}(\mathcal{X})$ and recall that, according to equation (5), we have that
@@
Let $x^{\prime}_{0}\in\mathcal{X}$ be such that $f(x_{0})=\mathbf{d}_{\mathcal{X}}(x_{0},x^{\prime}_{0})-f(x_{0}^{\prime})$. Notice that then, by equation (5), we have both
@@
Adding these two expressions together we obtain that
@@
$\inf_{x\in\mathcal{X}}\|f-\mathbf{d}_{\mathcal{X}}(x,\cdot)\|_{\infty}\leq\min\big{(}\|f-\mathbf{d}_{\mathcal{X}}(x_{0},\cdot)\|_{\infty},\|f-\mathbf{d}_{\mathcal{X}}(x_{0}^{\prime},\cdot)\|_{\infty}\big{)}<\frac{\delta}{2}.$
Since $f\in\operatorname{E}(\mathcal{X})$ was arbitrary, this proves that $\operatorname{hcdef}(\mathcal{X})<\frac{\delta}{2}$ from which the claim follows. ∎